# Rolled Gaussian process models for curves on manifolds

### v1

@@ 3.3 A parametric Gaussian process model for random curves on $\mathbb{R}^{d}$
where $\Phi=\left\{\phi(t_{1}),\ldots,\phi(t_{r})\right\}\in\mathbb{R}^{k\times r}$ and $Z=\left\{z(t_{1}),\ldots,z(t_{r})\right\}\in\mathbb{R}^{d\times r}$ are matrices.
4 Estimation and inference for curves in discrete time
4.1 Rolling, and the rolled model, in discrete time
In practice, curve data are recorded in discrete time, and a discrete version of the rolled Gaussian process model is accordingly needed. Let $t_{1},\ldots,t_{r}$ henceforth be times at which each curve is recorded and stored, with times $t_{i}=(i-1)/(r-1);i=1,\ldots,r$ assumed equally spaced on $[0,1]$, and with $r>k$. There is some loss of generality in assuming curves are recorded at common times, and that the observation times are equally spaced, but the former simplifies notation considerably, and both can be easily relaxed. We use upper case for the discrete-time version of the curve denoted in corresponding lower case; for example, $\Gamma:=\{\gamma(t_{1}),\ldots,\gamma(t_{r})\}\in M^{r}$ is curve $\gamma$ in discrete time.
The four maps in §3.1 have natural analogues for discrete curves: the unrolling, $\Gamma_{b}^{\downarrow}$, of $\Gamma$, is the discrete-time curve $\Gamma_{b}^{\downarrow}:=\{\gamma_{b}^{\downarrow}(t_{1}),\ldots,\gamma_{b}^{\downarrow}(t_{r})\}\in(T_{b}M)^{r}$, where $\gamma_{b}^{\downarrow}$ is the unrolling into $T_{b}M$ of the piecewise geodesic curve that has geodesic segments connecting consecutive points $\gamma(t_{i})$ and $\gamma(t_{i+1})$, such that $U^{\top}\Gamma_{b}^{\downarrow}\in\mathbb{R}^{d\times r}$ for a frame $U:\mathbb{R}^{d}\to T_{b}M$. Similarly, $(X)_{b}^{\downarrow{\Gamma}}$ with
as the $j$th component is the unwrapping of a discrete-time curve $X\in M^{r}$ with respect to the same piecewise geodesic $\gamma$ such that $U^{\top}(X)_{b}^{\downarrow{\Gamma}}\in\mathbb{R}^{d\times r}$. The unflattening maps, rolling and wrapping, for discrete curves follow in a similar way to these flattening ones; see Appendix 10.3.
For the discrete-time version of model $x\sim\mathcal{RGP}(m,K;b,U)$ under the parametric specification of $m$ and $K$ from §3.3, we write
@@ 4.1 Rolling, and the rolled model, in discrete time
Let $\Gamma=\{\gamma(t_{1}),\ldots,{\gamma}(t_{r})\}$ and $X=\{x(t_{1}),\ldots,{x}(t_{r})\}$ respectively be discrete-time versions of arbitrary curves ${\gamma}$ and $x$ in $M$. Fix $b\in M$ and frame $U$ for $T_{b}M$. When $x$ and $\gamma$ are outside the cut loci of each other, and $b$ is outside the cut locus of $\gamma(0)$, the unwrapping coordinates of $X$ are defined by the map
@@ 4.2 Estimation
Now consider the more challenging problem where the data are curves on $M$. For a sample of discrete-time curves $\{X_{i}\}_{i=1,\ldots,n}$, each $X_{i}\sim\mathcal{RMN}(M_{w},U_{w},V_{w};b,U)$ independently for a fixed point $b\in M$ and basis $U$ of $T_{b}M$, the goal is to estimate parameters $M_{w},U_{w},$ and $V_{w}$. Parameters $M_{w},V_{w},U_{w}$ may be estimated by maximising a likelihood written using (4.4) for the mean and covariance parameters in (4.3); moreover, $H(\Gamma;\Gamma)=M_{w}\Phi$.
Of course, $\Gamma$, is not known a priori; under the generative model it depends on $M_{w}$, which is to be estimated. Two candidate estimators are:
@@
related to (4.5), in which $\hat{M}_{w}\in\{\hat{M}^{\text{LS}}_{w},\hat{M}^{\text{MLE}}_{w}\}$. If $\hat{M}_{w}$ is taken to be $\hat{M}^{\text{LS}}_{w}$, then it can be evaluated once at the outset, followed by iterative evaluations of (4.8) and (4.9). If $\hat{M}_{w}$ is taken to be $\hat{M}^{\text{MLE}}_{w}$ then, owing to its dependence on $\hat{U}_{w}$ and $\hat{V}_{w}$, it becomes necessary to evaluate (4.7), (4.8) and (4.9) iteratively. One reason that $\hat{M}^{\text{MLE}}_{w}$ might be preferred in spite of this extra computational cost is that in the case where the manifold $M$ has non-positive sectional curvatures then $\hat{M}^{\text{MLE}}_{w}$ together with $\hat{U}_{w}$ $\hat{V}_{w}$ in (4.8, 4.9) are asymptotically, for large $n$, the maximum likelihood estimators, and are exactly so when $\Gamma$ is the data-generating curve.
Another estimator of $\Gamma$ is based on the discretization of the sample Fréchet mean curve, when it exists, which can then be used to estimate $M_{w}$. The discrete version of the sample Fréchet mean, defined in (2.3), is $\hat{\Gamma}_{\text{Fre}}=\{\hat{\gamma}_{\text{Fre}}(t_{1}),\ldots,\hat{\gamma}_{\text{Fre}}(t_{n})\}$. An estimator of $M_{w}$ based on the unwrapped coordinates with respect to $\hat{\Gamma}_{\text{Fre}}$ is
@@
In practice, especially for data on $M$ with small variability for each $t$, we have found barely any appreciable difference between estimators $\hat{M}^{\text{Fre}}_{w}$, $\hat{M}^{\text{LS}}_{w}$, $\hat{M}^{\text{MLE}}_{w}$. However, $\hat{M}_{w}^{\text{Fre}}$ has practical advantages in having a closed-form expression in terms of the sample Fréchet mean curve, $\hat{\Gamma}_{\text{Fre}}$, and having a natural interpretation as being expressible via the unrolling of $\hat{\Gamma}_{\text{Fre}}$ onto $T_{b}M$.
Proposition 4.2 (Estimator $\hat{M}_{w}^{\text{Fre}}$ via unrolling of the sample Fréchet mean).
Assume that the sample Fréchet mean curve in (2.3) exists and is unique. Then,
@@
Under the conditions of Theorem 3.3 which relates the Fréchet mean and the rolled mean, we prove that the estimator $\hat{M}^{\text{Fre}}_{w}$ is consistent for $M_{w}$ whenever the sample Fréchet mean curve, $\hat{\gamma}_{\text{Fre}}$ consistently estimates $\gamma_{\text{Fre}}$, under rolled Gaussian process model $x\sim\mathcal{RGP}(m,K;b,U)$ in a topology strong enough to ensure convergence of the parallel transport maps.
Theorem 4.2 (Consistency of estimator $\hat{M}^{\text{Fre}}_{w}$).
Assume that the population Fréchet mean curve $\gamma_{\text{Fre}}$ in (2.2) exists and is unique. Suppose $\hat{\gamma}_{\text{Fre}}$ converges in probability, as $n\to\infty$, in the $C^{1}$ topology to the population mean curve $\gamma$ under the rolled Gaussian process model. Then, under any of the conditions in Theorem 3.3, as $n\to\infty$, $\hat{M}^{\text{Fre}}_{w}$ converges in probability to $M_{w}$.
Remark 4.1.
$C^{1}$ convergence of curves in Theorem 4.2 is stronger than what is necessary to establish consistency of the estimator $\hat{M}^{\text{Fre}}_{w}$, since only necessary is convergence of parallel transport maps along their piecewise geodesic approximations at the finite times $t_{1},\ldots,t_{m}$.
@@ 4.3 A two-sample test for equality of mean curves
The flattening operation using the unrolling and unwrapping maps enable development of inferential tools for samples of curves on $M$. We briefly discuss how this can be done for a two-sample test of equality of means of two samples $\{X_{1}^{(1)},\ldots,X_{n_{1}}^{(1)}\}$ and $\{X_{1}^{(2)},\ldots,X_{n_{2}}^{(2)}\}$ of curves on $M$, which we anticipate will be relevant in many applications. A resampling-based test for equality of mean curves, using a test statistic analogous to the Hotelling $T^{2}$ statistic in classical multivariate setting, is as follows. Suppose that
@@
in which for $g=1,2$ the mean parameters are possibly different, but the covariance parameters are assumed common. In this context, the hypothesis to test is $H_{0}:M_{w}^{(1)}=M_{w}^{(2)}$ versus the alternative $H_{1}:M_{w}^{(1)}\neq M_{w}^{(2)}$. From §4.2, given estimates $\hat{M}^{(g)}_{w},\hat{U}^{(g)}_{w},\hat{V}^{(g)}_{w}$ for each sample, the analogue of the Hotelling $T^{2}$ statistic is
@@ 5 Practical calculations in embedded coordinates
rolling $\gamma_{b}^{\uparrow}$ is defined as previously as the unique inverse of unrolling, and the unwrapping and wrapping maps are as defined before in (3.5-3.6) with elements of $M$ or $T_{b}M$ being vectors in $\mathbb{R}^{q}$, and each parallel transport being a matrix, $\mathbb{R}^{q\times q}$. For the rolled Gaussian process $x\sim\mathcal{RGP}(m,K;b,U)$ in embedded coordinates, basis $U=(u_{1},\ldots,u_{d})\in\mathbb{R}^{q\times d}$ is a matrix with $d$ orthonormal columns, and the Gaussian process $z\sim\mathcal{GP}(m,K)$ is such that $z(t)=\sum_{i=1}^{d}e_{i}z(t)$ where $\{e_{i}\}$ is the standard basis for $\mathbb{R}^{d}$; then for $\tilde{m}_{b}:\tilde{m}_{b}(t)=b+\sum_{i=1}^{d}u_{i}m(t)$ and $y:y(t)=b+\sum_{i=1}^{d}u_{i}z(t)$, $\gamma=\tilde{m}_{b}^{\uparrow}$ is the rolled mean, and finally $x=y_{b}^{\uparrow\gamma}$.
Recall that the discrete-time model (4.1) for curves on $M$ is based upon rolling and wrapping curves from the underlying model (3.8) for curves on $\mathbb{R}^{d}$. Let $\tilde{M}=b1_{r}^{\top}+UM_{w}\Phi$ be the mean curve in (3.8) identified as a curve in $T_{b}M$, and $\Gamma=(\tilde{M})^{\uparrow}_{b}$ be its rolling onto $M$; for $Z$ a random curve from (3.8), $Y=b1_{r}^{\top}+UZ$ is the corresponding curve in $T_{b}M$, and $X=Y_{b}^{\uparrow\Gamma}$ its wrapping with respect to $\Gamma$ onto $M$.
@@ 6.1 A prescribed model for heteroscedastic curves on $\mathbb{S}^{2}$
Here we give an example for $M=\mathbb{S}^{2}$, of the steps described in the preceding section to simulate from (4.1) for curves on $\mathbb{S}^{2}$: first define a model (3.8) in $\mathbb{R}^{2}$, with prescribed parameters $M_{w},U_{w},V_{w}$, simulate for the model in $\mathbb{R}^{2}$ then roll and wrap to produce curves $\mathbb{S}^{2}$ that are samples from (4.1). Let $s_{1},\ldots,s_{k}$ be an equally spaced sequence of points spanning $[0,1]$, $M_{w}\in\mathbb{R}^{2\times k}$ be the matrix with $i$th column equal to $3/4\cdot\left[(1,1)+(1/2+s_{i}^{2}/2)\cdot\left\{\cos(5s_{i}),\sin(5s_{i})\right\}\right]^{\top}$, and $k=10$. The mean curve, $m(t)=M_{w}\phi(t)$, evaluated discretely as $M_{w}\Phi$ at $r=100$ values of $t$ equally spaced on $[0,1]$, is shown as the red line in Fig. 3(a). To prescribe the covariance parameters, let $U_{w}=I_{2}$, and let $V_{w}$ be the matrix with $(i,j)$th element equal to $a_{i}a_{j}\rho^{|i-j|}$, with $a_{i}=1+3/4\cdot\cos(2\pi i/k)$ and $\rho=0.9$. This prescription produces curves in $\mathbb{R}^{2}$ with smooth variation and some moderate heteroscedasticity. A random curve from this model is $z_{i}(t)=W_{i}\phi(t)$, with $W_{i}\sim\mathcal{MN}(M_{w},U_{w},V_{w})$, evaluated numerically as $Z_{i}=W_{i}\Phi$. Fig. 3(a) shows a sample of $n=10$ curves from this model.
@@ 6.2 Curves on $\operatorname{Sym}_{>0}(2)$
We equip the three-dimensional manifold $\operatorname{Sym}_{>0}(2)$ with the affine-invariant metric, $d(\cdot,\cdot)$, which makes it negatively curved (e.g. Pennec et al., 2019, Chapter 3); see 10.5 for definition of $d(\cdot,\cdot)$. The global coordinates of $\mathbb{R}^{3}$ is used to define a model with parameters $M_{w},U_{w},V_{w}$ and roll onto $\operatorname{Sym}_{>0}(2)$, then from simulated data try to infer the parameters. With $s_{1},\ldots,s_{k}$ an equally spaced sequence of points spanning $[0,1]$ and $k=5$, let $M_{w}\in\mathbb{R}^{3\times k}$ be the matrix with $i$th column equal to $3/4\cdot\left[(1/5)\cdot\left\{\cos(5s_{i}),\sin(5s_{i}),s_{i}\right\}\right]^{\top}$, $U_{w}=I_{3}$, and $V_{w}$ be the matrix with $(i,j)$th element equal to $10^{-}3a_{i}a_{j}\rho^{|i-j|}$, with $a_{i}$, $\rho$ as in §6.1. We take the tangent point $b=I_{2}$ and basis for tangent space $U=I_{3}$, then simulate data $\{X_{i}\}$ as in §6.1. Table 1 shows numerical results from simulating the data then fitting the model to the simulated data, to investigate convergence of estimators $\hat{M}_{w}=\hat{M}_{w}^{\text{Fre}}$ from (4.10) and $\hat{U}_{w}$ and $\hat{V}_{w}$ from (4.8) and (4.9) to their data-generating values. Results in the table show each estimator, with respect to the indicated metric, getting closer to towards its population counterpart as the sample size, $n$, increases; a Mahalaonobis-type metric is used for $\hat{M}_{w}$ and $d(\cdot,\cdot)$ is used for $\hat{U}_{w}$ and $\hat{V}_{w}$.

### v2

@@ 6 Estimation and testing for discretely observed curves under a parametric model
where $\Phi=\left\{\phi(t_{1}),\ldots,\phi(t_{r})\right\}\in\mathbb{R}^{k\times r}$ and $Z=\left\{z(t_{1}),\ldots,z(t_{r})\right\}\in\mathbb{R}^{d\times r}$ are matrices. Then, the corresponding discrete-time version of model $x\sim\mathcal{RGP}(m,K;b,U)$ is denoted by
@@
Given curves $\gamma$ and $x$ in $M$, consider, respectively, their discrete-time versions $\Gamma=\{\gamma(t_{1}),\ldots,{\gamma}(t_{r})\}$ and $X=\{x(t_{1}),\ldots,{x}(t_{r})\}$. Given $(b,U)$, when $x$ and $\gamma$ are outside the cut loci of each other, and $b$ is outside the cut locus of $\gamma(0)$, the unwrapping coordinates of $X$ are defined by the map
@@ 6.1 Estimation
We now return to our setting consisting of a sample of discrete-time curves $\{X_{i}\}_{i=1,\ldots,n}$, where each $X_{i}\overset{ind}{\sim}\mathcal{RMN}(M_{w},U_{w},V_{w};b,U)$ for a fixed point $b\in M$ and basis $U$ of $T_{b}M$. We present multiple estimators of $M_{w}$, each of which results in different estimators of $U_{w}$ and $V_{w}$. The first set of estimates are obtained by maximising a likelihood written using (6.5) for the mean and covariance parameters in (6.4); moreover, $H(\Gamma;\Gamma)=M_{w}\Phi$. Alternatively, consider the two estimators:
@@
related to (6.6), in which $\hat{M}_{w}\in\{\hat{M}^{\text{LS}}_{w},\hat{M}^{\text{MLE}}_{w}\}$. If $\hat{M}_{w}$ is taken to be $\hat{M}^{\text{LS}}_{w}$, then it can be evaluated once at the outset, followed by iterative evaluations of (6.9) and (6.10). If $\hat{M}_{w}$ is taken to be $\hat{M}^{\text{MLE}}_{w}$ then, owing to its dependence on $\hat{U}_{w}$ and $\hat{V}_{w}$, it becomes necessary to evaluate (6.8), (6.9) and (6.10) iteratively. One reason that $\hat{M}^{\text{MLE}}_{w}$ might be preferred in spite of this extra computational cost is that in the case where the manifold $M$ has non-positive sectional curvatures then $\hat{M}^{\text{MLE}}_{w}$ together with $\hat{U}_{w}$ $\hat{V}_{w}$ in (6.9, 6.10) are asymptotically, for large $n$, the maximum likelihood estimators, and are exactly so when $\Gamma$ is the data-generating curve.
Another estimator of $\Gamma$ is based on the discretization of the sample Fréchet mean curve, when it exists, which can then be used to estimate $M_{w}$. This is closely related to the discretization of the estimator $\hat{m}_{r}$ in (5.1). Let $\hat{\Gamma}_{\text{Fre}}=\{\hat{\gamma}_{\text{Fre}}(t_{1}),\ldots,\hat{\gamma}_{\text{Fre}}(t_{n})\}$ be discrete version of the sample Fréchet mean $\hat{\gamma}_{\text{Fre}}$. An estimator of $M_{w}$ based on the unwrapped coordinates with respect to $\hat{\Gamma}_{\text{Fre}}$ is
@@
The estimator $\hat{M}^{\text{Fre}}_{w}$ is exactly the one described in Remark 5.1 as the projection of the mean of the unwrapped discrete curves $X_{i}$ onto the $k$-dimensional subspace spanned by $\Phi$, and has a closed-form expression.
Proposition 6.2 (Estimator $\hat{M}_{w}^{\text{Fre}}$ via unrolling of $\hat{\gamma}_{\text{Fre}}$).
Under assumption A1
@@
As discussed in Remark 5.1, under assumptions A1-A5, we have $\|\hat{M}^{\text{Fre}}_{w}-M_{w}\|_{F}=O_{P}(n^{-1/2})$. In practice, especially for data on $M$ with small variability for each $t$, we found barely any appreciable difference between estimators $\hat{M}^{\text{Fre}}_{w}$, $\hat{M}^{\text{LS}}_{w}$, $\hat{M}^{\text{MLE}}_{w}$.
@@ 6.2 A two-sample test for equality of mean curves
The flattening operation using the unrolling and unwrapping maps enable development of inferential tools for samples of curves on $M$. As a simple illustration, we briefly discuss how this can be done for a two-sample test of equality of means of two samples $\{X_{1}^{(1)},\ldots,X_{n_{1}}^{(1)}\}$ and $\{X_{1}^{(2)},\ldots,X_{n_{2}}^{(2)}\}$ of curves on $M$, which we anticipate will be relevant in many applications.
When $n_{1}+n_{2}$ is large enough, it is sensible to consider a test statistic similar to the Hotelling’s $T^{2}$ statistic. This is so, since, under the rolled Gaussian process model, the unwrapping coordinates for the two samples $\{H(X_{1},\hat{\Gamma}_{\text{fre}})^{(1)},\ldots,H(X_{n_{1}},\hat{\Gamma}_{\text{fre}})^{(1)}\}$ and $\{H(X_{1},\hat{\Gamma}_{\text{fre}})^{(2)},\ldots,H(X_{n_{2}},\hat{\Gamma}_{\text{fre}})^{(2)}\}$ are approximately Gaussian distributed, where $\hat{\Gamma}_{\text{fre}}$ is the discretized sample Fréchet mean curve of the pooled sample. If the population version $\Gamma_{\text{fre}}$ of the pooled sample is used in place of $\hat{\Gamma}_{\text{fre}}$, then the unwrapping coordinates are exactly Gaussian distributed since they are finite-dimensional projections of the covariant Gaussian vector in Proposition 4.1.
Now, suppose
@@
in which for $g=1,2$ the mean parameters are possibly different, but the covariance parameters are assumed common. We wish to test $H_{0}:M_{w}^{(1)}=M_{w}^{(2)}$ versus the alternative $H_{1}:M_{w}^{(1)}\neq M_{w}^{(2)}$. From §6.1, given estimates $\hat{M}^{(g)}_{w},\hat{U}^{(g)}_{w},\hat{V}^{(g)}_{w}$ for each sample, the analogue of the Hotelling $T^{2}$ statistic is
@@
The two-sample test offers one compelling illustration of the inferential utility of the rolled Gaussian process model for curves observed in discrete time. Evidently, the estimators of $M_{w},U_{w},V_{w}$ may be used to develop many other inferential procedures. We also note that various multivariate test statistics that testing equality of the mean may be used; our purpose in this section is to merely demonstrate the utility of one such sensible one.
@@ 7 Practical calculations in embedded coordinates
rolling $\gamma_{b}^{\uparrow}$ is defined as previously as the unique inverse of unrolling, and the unwrapping and wrapping maps are as defined before in (3.5-3.6) with elements of $M$ or $T_{b}M$ being vectors in $\mathbb{R}^{q}$, and each parallel transport being a matrix, $\mathbb{R}^{q\times q}$. For the rolled Gaussian process $x\sim\mathcal{RGP}(m,K;b,U)$ in embedded coordinates, basis $U=(u_{1},\ldots,u_{d})\in\mathbb{R}^{q\times d}$ is a matrix with $d$ orthonormal columns, and the Gaussian process $z\sim\mathcal{GP}(m,K)$ is such that $z(t)=\sum_{i=1}^{d}e_{i}z(t)$ where $\{e_{i}\}$ is the standard basis for $\mathbb{R}^{d}$; then for $m_{b}(t)=b+\sum_{i=1}^{d}u_{i}m(t)$ and $y:y(t)=b+\sum_{i=1}^{d}u_{i}z(t)$, $\gamma={m}_{b}^{\uparrow}$ is the rolled mean, and finally $x=y_{b}^{\uparrow\gamma}$.
Recall that the discrete-time model (6.2) for curves on $M$ is based upon rolling and wrapping curves from the underlying model (6.1) for curves on $\mathbb{R}^{d}$. Let $M_{b}=b1_{r}^{\top}+UM_{w}\Phi$ be the mean curve in (6.1) identified as a curve in $T_{b}M$, and $\Gamma=(M_{b})^{\uparrow}_{b}$ be its rolling onto $M$; for $Z$ a random curve from (6.1), $Y=b1_{r}^{\top}+UZ$ is the corresponding curve in $T_{b}M$, and $X=Y_{b}^{\uparrow\Gamma}$ its wrapping with respect to $\Gamma$ onto $M$.
@@ 8.1 A prescribed model for heteroscedastic curves on $\mathbb{S}^{2}$
Here we give an example for $M=\mathbb{S}^{2}$, of the steps described in the preceding section to simulate from (6.2) for curves on $\mathbb{S}^{2}$: first define a model (6.1) in $\mathbb{R}^{2}$, with prescribed parameters $M_{w},U_{w},V_{w}$; simulate from the model in $\mathbb{R}^{2}$; then roll and wrap to produce curves $\mathbb{S}^{2}$ that are samples from (6.2). Let $s_{1},\ldots,s_{k}$ be an equally spaced sequence of points spanning $[0,1]$, $M_{w}\in\mathbb{R}^{2\times k}$ be the matrix with $i$th column equal to $3/4\cdot\left[(1,1)+(1/2+s_{i}^{2}/2)\cdot\left\{\cos(5s_{i}),\sin(5s_{i})\right\}\right]^{\top}$, and $k=10$. The mean curve, $m(t)=M_{w}\phi(t)$, evaluated discretely as $M_{w}\Phi$ at $r=100$ values of $t$ equally spaced on $[0,1]$, is shown as the red line in Fig. 3(a). To prescribe the covariance parameters, let $U_{w}=I_{2}$, and let $V_{w}$ be the matrix with $(i,j)$th element equal to $a_{i}a_{j}\rho^{|i-j|}$, with $a_{i}=1+3/4\cdot\cos(2\pi i/k)$ and $\rho=0.9$. This prescription produces curves in $\mathbb{R}^{2}$ with smooth variation and some moderate heteroscedasticity. A random curve from this model is $z_{i}(t)=W_{i}\phi(t)$, with $W_{i}\sim\mathcal{MN}(M_{w},U_{w},V_{w})$, evaluated numerically as $Z_{i}=W_{i}\Phi$. Fig. 3(a) shows a sample of $n=10$ curves from this model.
@@ 8.2 Curves on $\operatorname{Sym}_{>0}(2)$
We equip the three-dimensional manifold $\operatorname{Sym}_{>0}(2)$ with the affine-invariant metric, $d(\cdot,\cdot)$, which makes it negatively curved (e.g. Pennec et al., 2019, Chapter 3); see B.3 for definition of $d(\cdot,\cdot)$. The global coordinates of $\mathbb{R}^{3}$ is used to define a model with parameters $M_{w},U_{w},V_{w}$ and roll onto $\operatorname{Sym}_{>0}(2)$, and then infer the parameters from simulated data. With $s_{1},\ldots,s_{k}$ an equally spaced sequence of points spanning $[0,1]$ and $k=5$, let $M_{w}\in\mathbb{R}^{3\times k}$ be the matrix with $i$th column equal to $3/4\cdot\left[(1/5)\cdot\left\{\cos(5s_{i}),\sin(5s_{i}),s_{i}\right\}\right]^{\top}$, $U_{w}=I_{3}$, and $V_{w}$ be the matrix with $(i,j)$th element equal to $10^{-}3a_{i}a_{j}\rho^{|i-j|}$, with $a_{i}$, $\rho$ as in §8.1. We take the tangent point $b=I_{2}$ and basis for tangent space $U=I_{3}$, then simulate data $\{X_{i}\}$ as in §8.1. Table 1 shows numerical results from simulating the data then fitting the model to the simulated data, to investigate convergence of estimators $\hat{M}_{w}=\hat{M}_{w}^{\text{Fre}}$ from (6.11) and $\hat{U}_{w}$ and $\hat{V}_{w}$ from (6.9) and (6.10) to their data-generating values. Results in the table show each estimator, with respect to the indicated metric, getting closer to towards its population counterpart as the sample size, $n$, increases; a Mahalaonobis-type metric is used for $\hat{M}_{w}$ and $d(\cdot,\cdot)$ is used for $\hat{U}_{w}$ and $\hat{V}_{w}$.